OpenStudy (anonymous):

estimate the area of the circle to the nearest unit. A- 6,644 cm^2 B- 72 cm^2 C- 144 cm^2 D- 1,661 cm^2

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

For a circle \[\large A = \pi r^2\] where 'r' is the radius Now we are given the Diameter....which is 2 times the radius... So the radius is just that 46/2 = 23 So now we have \[\large A = \pi (23)^2\] \[\large A = 529\pi\] \[\large A = 1661cm^2\]

OpenStudy (anonymous):

thanks!

OpenStudy (johnweldon1993):

Anytime :)